OpenStudy (anonymous):

A point T on a segment with endpoints D(1,4) and F (7,1) partitions the segment in a 2:1 ratio. find T.

OpenStudy (anonymous):

please help medal will be given

OpenStudy (anonymous):

@zepdrix can you help me with this

OpenStudy (jdoe0001):

ahemm thought you've covered this already =)

OpenStudy (jdoe0001):

like quite a few times by now

OpenStudy (anonymous):

i have short term memory and i am very bad at math :-)

OpenStudy (jdoe0001):

hmmm maybe it wasn't ... you I recall anyhow.. gimme a sec

OpenStudy (anonymous):

SOMEONE HELPPP

OpenStudy (jdoe0001):

 \(\bf D(1,4)\qquad F(7,1)\qquad ratio1=2\qquad ratio2=1\qquad 2:1\\ \quad \\ \quad \\ \cfrac{DT}{FT}=\cfrac{ratio1}{ratio2}\implies ratio2\cdot DT=ratio1\cdot FT\quad \textit{dividing by T}\\ \quad \\ ratio2\cdot D=ratio1\cdot F\implies 1(1,4)=2(7,1)\\ \quad \\\qquad {\color{blue}{ T=\left(\cfrac{\textit{sum of "x" values}}{ratio1+ratio2}\quad ,\quad \cfrac{\textit{sum of "y" values}}{ratio1+ratio2}\right)}}\\ \quad \\ \qquad thus\qquad \\ \quad \\ T=\left(\cfrac{(1\cdot 1)+(2\cdot 7)}{2+1}\quad ,\quad \cfrac{(1\cdot 4)+(2\cdot 1)}{2+1}\right)\)

OpenStudy (anonymous):

THANK YOU SO MUCH. i wont forget this time :)

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

yw